OpenStudy (annie96):

1. Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y (Points : 1) The solution is (1, 5) The solution is (21, 0) There are an infinite number of solutions. There is no solution. Question 2. Solve, using the substitution method. y = 2x + 6 3x – y = 6 (Points : 1) (14, 12) (12, 30) There is no solution. There are an infinite number of solutions.

OpenStudy (yanasidlinskiy):

Do you have any clue?? On how to do this? If not, i can show you for number 2.

OpenStudy (annie96):

ok

OpenStudy (yanasidlinskiy):

Since you know what y= already, plug it in for the second equation which gives you: 3x-(2x+6) = 6. Now you have to distribute the negative. 3x-2x-6=6. So far with me?

OpenStudy (yanasidlinskiy):

Your next step is to combine like terms on both sides. 3x-2x -----Those are like terms which give you: 1x-6=6. To get x by itself you have to do the opposite of what is added or subtracted to the x. So you would add the (6) to both sides....(Remember it is with the x.) 1x=12 -----Now divide by 1 to get x. x=12!

OpenStudy (annie96):

the answer would be A.

OpenStudy (yanasidlinskiy):

Lets find out! We are not done yet. We solved for x already. Now since we know what x is, we can plug in x into either of the equations to get y. Y=2x+6. Now we plug in the 12 for x. y=2(12) +6 y=24+6 y=30

OpenStudy (yanasidlinskiy):

(12,30)

OpenStudy (yanasidlinskiy):

Got it?